OpenStudy (anonymous):

If radical a / radical b = x, which statement must be true? a) x^2 = a/b b) radical x = a/b c) |x| a/b d) x = a/b

jimthompson5910 (jim_thompson5910):

\[\Large \frac{\sqrt{a}}{\sqrt{b}} = x\] right?

OpenStudy (anonymous):

Yes!

jimthompson5910 (jim_thompson5910):

We have square roots on the left side. So one natural thing to do is to square both sides to get rid of them. \[\Large \frac{\sqrt{a}}{\sqrt{b}} = x\] \[\Large \left(\frac{\sqrt{a}}{\sqrt{b}}\right)^2 = x^2\] \[\Large \frac{(\sqrt{a})^2}{(\sqrt{b})^2} = x^2\] \[\Large \frac{a}{b} = x^2\] \[\Large x^2 = \frac{a}{b}\]

OpenStudy (anonymous):

I literally just got what the question was asking me after five times of seeing it. Thank you so, so, so much. :)

jimthompson5910 (jim_thompson5910):

you're welcome and keep practicing